Sally bought 3 choclate bars and a pack of gum and paid $1.75. Jake bought 2 chocolate bars and 4 packs of gum and paid $2.00. Write a system of algebraic equations using c=chocolate bar, and g=gum to represent the given information. Solve using the adding method or substitution to fing the cost of a chocolate bar and the cost of a pack of gum.
Answer by stanbon(75887)

Sally bought 3 choclate bars and a pack of gum and paid $1.75. Jake bought 2 chocolate bars and 4 packs of gum and paid $2.00. Write a system of algebraic equations using c=chocolate bar, and g=gum to represent the given information. Solve using the adding method or substitution to find the cost of a chocolate bar and the cost of a pack of gum.
------------
EQUATIONS:
3c+g = 175 cents
2c+4g = 200 cents
-----------------
Multiply the 1st equation by 4 to get:
12c + 4g = 700
2c+4g = 200
---------------
Subtract the 2nd equation from the 1st to get:
10c = 500
c = 50 cents (cost of a chocolate bar)
----------
Substitute into 3c+g = 175 to solve for g:
3*50 + g= 175
g = 25 cents (cost of a pack of gum)
==============
Cheers,
Stan H.